OpenStudy (anonymous):

How do I find x in this problem? I am thinking it is going to require a change of base? 3^x=(1/3)^(x+1)

OpenStudy (anonymous):

Recall that \(\dfrac{1}{3}=3^{-1}\), so \[3^x=\frac{1}{3^{x+1}}~~\iff~~3^x=3^{-(x+1)}\]

OpenStudy (anonymous):

awsome thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I was thinking it was along those lines.